### v1

# Evolving Hard Maximum Cut Instances for Quantum Approximate Optimization Algorithms

Abstract.
Variational quantum algorithms, such as the Recursive Quantum Approximate Optimization Algorithm (RQAOA), have become increasingly popular, offering promising avenues for employing Noisy Intermediate-Scale Quantum devices to address challenging combinatorial optimization tasks like the maximum cut problem. In this study, we utilize an evolutionary algorithm equipped with a unique fitness function. This approach targets hard maximum cut instances within the latent space of a Graph Autoencoder, identifying those that pose significant challenges or are particularly tractable for RQAOA, in contrast to the classic Goemans and Williamson algorithm. Our findings not only delineate the distinct capabilities and limitations of each algorithm but also expand our understanding of RQAOA’s operational limits. Furthermore, the diverse set of graphs we have generated serves as a crucial benchmarking asset, emphasizing the need for more advanced algorithms to tackle combinatorial optimization challenges. Additionally, our results pave the way for new avenues in graph generation research, offering exciting opportunities for future explorations.
variational quantum algorithms, RQAOA, Goemans and Williamson algorithm, graph autoencoder
conference: Genetic and Evolutionary Computation Conference (GECCO ’25); July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25), July 14–18, 2025, Malaga, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXX
1. Introduction
In the era of noisy intermediate-scale quantum (NISQ) technology, the practical use of quantum computing is being increasingly pursued, despite challenges like the scarcity of quantum resources and the noise in quantum gates (Preskill, 2018). Key research areas such as combinatorial optimization (Farhi et al., 2014), quantum machine learning (Benedetti et al., 2019), and quantum chemistry (Moll et al., 2018) are recognized as potential fields where NISQ devices might surpass best classical methods, thereby showcasing a quantum advantage.
The Recursive Quantum Approximate Optimization Algorithm (RQAOA) (Bravyi et al., 2020) was developed to improve upon the Quantum Approximate Optimization Algorithm (QAOA) for combinatorial optimization problems like maximum cut. While QAOA (Farhi et al., 2014) was initially introduced as a quantum heuristic, research indicates it struggles to surpass the Goemans and Williamson (GW) algorithm (Goemans and Williamson, 1995), a classical method with strong approximation guarantees, at constant depth (Hastings, 2019; Farhi et al., 2020; Marwaha, 2021; Barak and Marwaha, 2022). In contrast, RQAOA employs a recursive strategy that iteratively refines solutions, demonstrating superior performance both numerically and analytically (Bravyi et al., 2021, 2022; Bae and Lee, 2024). This makes RQAOA a promising advancement in quantum optimization.
Research has shown that certain algorithms underperform in specific scenarios, often labeled as hard or worst-case instances (Ko et al., 1986). Yet, guidance on how to generate or identify these challenging instances of studies is typically lacking in the literature. For instance, the RQAOA paper highlights an infinite family of $d$-regular bipartite graphs where the QAOA lags behind the GW algorithm when the depth (number of layers) is $O(log\textit{n})$, where n is the number of nodes (Bravyi et al., 2020). However, the paper does not provide instructions for constructing these graphs. In our work, we seek to bridge this gap by generating instances that underscore the differences between quantum algorithms and classical optimization algorithms in tackling the maximum cut problem, with a particular focus on the RQAOA and GW algorithms. While various heuristics have been explored in (Dunning et al., 2018), we selected the GW algorithm due to its well-established approximation bounds and extensive documentation of both its strengths and limitations, making it a suitable benchmark for comparison. We employ an evolutionary method using the classical Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen and Ostermeier, 2001; Hansen et al., 2003) paired with a fitness function specifically tailored for this purpose. This strategy generates a variety of instances in the latent space of a thoroughly trained Graph Autoencoder (GAE). These instances are designed to be easily solvable by one algorithm while posing substantial challenges to the other, thereby providing a clearer comparison of their relative strengths and weaknesses.
We deliver several significant contributions:
  • We have developed a method that effectively pinpoints maximum cut instances that either pose significant challenges or are considerably more manageable for the RQAOA in comparison to the GW algorithm.
  • Through an extensive analysis of these instances across various sizes of the maximum cut problem, our generated instances serve as valuable benchmarks for evaluating and contrasting the effectiveness of different algorithms in addressing maximum cut challenges. This benchmarking not only underscores performance discrepancies among algorithms but also aids in advancing quantum computing methodologies.
  • Our research brings a fresh perspective to graph generation studies within the network science community, spurring further investigation and innovation in this domain.
2. PRELIMINARIES
Maximum cut
The maximum cut problem in graphs is one of the classical and well-studied NP-hard combinatorial optimization problems. Given an undirected graph $G=(V,E)$, the goal is to compute a partitioning into two sets $S\subseteq V$ and $V\setminus S$ such that the number of edges
$C(S)=|\{e\in E\mid e\cap S\not=\emptyset\wedge e\cap(V\setminus S)\not=\emptyset\}|$
crossing the two sets is maximal. The maximum cut problem is APX-hard (Papadimitriou and Yannakakis, 1991) and can be approximated using the GW algorithm (Goemans and Williamson, 1995). Specifically, this approach involves solving a relaxed version of the original problem. We first tackle an optimization problem in the continuous domain $[-1,1]^{N}$ using a semidefinite program, which provides us with a real-valued solution associated with a cost denoted as $C_{\text{relaxed}}$. Then, a sampling process, guided by probabilities derived from the entries of the real-valued solution (via a random projection method), is used to produce a feasible solution in the form of a binary string. It is important to note that the cost of the generated binary string is guaranteed to be at most equal to $C_{\text{relaxed}}$, and it is possible that $C_{\text{max}}<C_{\text{relaxed}}$. However, the expected value obtained by the cost function using this method is proven to be less than $\alpha C_{\text{max}}$, where $\alpha\approx 0.878$, for any input graph. In other words, this is an $\alpha$-approximation algorithm.
The maximum cut problem has received significant attention in the area of quantum optimization and different quantum algorithms have been designed. As discussed earlier, QAOA faces significant limitations due to its intrinsic locality and symmetry constraints, with substantial theoretical research having established no-go theorems in this area. To address these challenges, RQAOA (Bravyi et al., 2020) was introduced as an alternative quantum optimization algorithm specifically for combinatorial optimization problems. During each iteration of RQAOA, the problem size—represented as a graph or hypergraph—is reduced by one or more elements, based on correlations between variables derived from QAOA. This reduction process introduces non-local effects by creating new connections between nodes that were not previously linked, thus circumventing the locality constraints of QAOA. Finally, the iterative process stops once the number of variables is below a certain predefined threshold $n_{c}$ ($n_{c}\approx O(1)<<|V|$) such that a classical solver can be used to solve the remaining graph. In this study, we focus on depth-1 RQAOA, which can be classically simulated with a runtime of $O(n^{4})$ (Patel et al., 2024). Understanding the performance of these algorithms is challenging and pointing out instances where quantum approaches perform significantly better or worse than the classical GW algorithm provides an important understanding of their performance in comparison to state-of-the-art classical approximation algorithms.
Figure 1 presents the early performance of the RQAOA ($n_{c}$ is set to 10) over the first four iterations on a specifically generated 20-node maximum cut instance where it outperforms the GW algorithm. The diagram details the expectation values for each edge during the RQAOA process. In every iteration, the algorithm selects the edge that exhibits the highest absolute expectation value for operation. In cases where several edges possess identical maximum values, one edge is chosen at random. This sequential representation clearly illustrates the strategic process RQAOA employs to make decisions at each step.
Figure 1. Visualization of RQAOA’s performance ($n_{c}=10$) over the first four iterations on a 20-node maximum cut instance, where it outperforms the GW algorithm. Blue dots represent nodes, grey lines indicate edges, and the orange-highlighted node and edge denote the selected operation. The diagram tracks expectation values for each edge at every iteration, with RQAOA choosing the edge with the highest absolute expectation value, selecting randomly in case of a tie.
[…]
Visualization of RQAOA performance ($n_{c}$ is set to 10) over the first four iterations for a 20-node maximum cut instance, where it outperforms the GW algorithm.
Graph Neural Networks (GNNs)
GNNs have been proposed to tackle learning problems involving non-Euclidean data (Ju et al., 2024; Zhu et al., 2022), among which graph autoencoders (GAEs) are successfully applied to unsupervised tasks to learn graph topologies (Kipf and Welling, 2016). GAEs encode vertices or the entire graph into a latent space, where several different encoding/decoding methods are devised. GAEs have been used to extract network representations and learn the empirical distributions thereof (Wu et al., 2020). Due to the famous graph isomorphism issues, the generating/decoding component of GNNs is non-trivial: several works target this issue, including GraphRNN (You et al., 2018) employing a sequential method to construct nodes and edges by learning from a representative set of graphs; GraphVAE (Simonovsky and Komodakis, 2018) that uses variational autoencoders (Kingma and Welling, 2014) to generate graphs from continuous embeddings.
Covariance Matrix Adaptation Evolution Strategy (CMA-ES)
CMA-ES (Hansen, 2006, 2016) is the state-of-the-art numerical algorithm for black-box optimization problems. CMA-ES employs a set/population of search points, which are sampled from a multivariate Gaussian distribution. CMA-ES selects the subset of points with better objective values and uses their information to update the mean and covariance of the Gaussian. CMA-ES has been successfully applied to various challenging applications and is known for its fast convergence and robustness (Hansen et al., 2009; Loshchilov and Hutter, 2016; Bonet-Monroig et al., 2023; Salimans et al., 2017).
3. Related work
In optimization, generating diverse problem instances is essential for effective benchmarking. The choice of instances significantly impacts algorithm evaluation, making it crucial to update benchmarks regularly to prevent overfitting specific cases. Assessing benchmark quality is challenging, as measuring diversity and representativeness is not straightforward. Typically, evaluation considers key dimensions such as feature space, performance space, and instance space (Bartz-Beielstein et al., 2020).
The literature highlights various methods for generating maximum cut instances used in benchmarking. A notable example is the Balasundarm-Butenko problem generator (Balasundaram and Butenko, 2005), which utilizes a framework designed for creating test functions within global optimization. This involves applying continuous formulations to combinatorial optimization problems. Another key tool is the machine-independent graph generator known as Rudy, introduced by (Helmberg and Rendl, 2000) which has been widely used in subsequent research to evaluate algorithms (Benson et al., 1999; Burer and Monteiro, 2001; Burer et al., 2002). In 2018, a comprehensive library (Dunning et al., 2018) containing 3,296 diverse instances was introduced, aggregating data from multiple sources, and further enriching the resources available for algorithm testing.
Evolutionary algorithms (EAs) have proven effective in generating diverse problem instances across various domains. Bossek et al. (2019) pioneered an EA-based approach for crafting the Travelling Salesperson Problem (TSP) instances using novel mutation operations. Neumann et al. (2019) developed a method for evolving diverse TSP instances and images using multi-objective indicators. Gao et al. (2021) introduced an EA-driven framework to generate problem instances classified as easy or hard based on distinct features. Addressing diversity in high-dimensional feature spaces, Marrero et al. (2023) proposed a Knapsack problem instance generator combining Principal Component Analysis with novelty search. More recently, research has explored evolving instances for the chance-constrained maximum coverage problem (Sadeghi Ahouei et al., 2024). This study departs from traditional approaches that rely on approximation ratios in the $(1+1)$ EA, instead incorporating performance ratio variances for improved evaluation.
4. Evolving Hard Maximum Cut Instances
GNNs have proven effective for generating graphs with diverse attributes. In our research, we use tools like the GAE to explore and identify specific graph structures. Our goal is to identify hard maximum cut instances. We begin by training a GAE with graph generators that are rooted in network science. After the GAE is adequately trained, we proceed to pinpoint particular graph instances within its complex latent space. The identification process is facilitated by the classical CMA-ES algorithm. This strategy employs a fitness function to assess scenarios where the RQAOA either surpasses or falls short compared to the GW algorithm.
Problem formulation
We aim to propose a method for evolving hard instances of the maximum cut problem, to challenge a recently proposed quantum search algorithm - RQAOA, against a well-studied classical algorithm - GW. We measure the largest number of cuts found by each algorithm on a graph instance $I$, denoted as $\textit{p}_{\text{RQAOA}}(\text{I})$, or $\textit{p}_{\text{GW}}(\text{I})$.
Graph instance representation
We choose to base our research on Permutation-Invariant Variational Autoencoder (PIGVAE) (Winter et al., 2021a) as it handles the graph isomorphism problem explicitly. The core principle of PIGVAE is that identical graphs should yield the same representation, regardless of the order of their nodes. It combines a variational autoencoder with a permuter model. This permuter model, trained in conjunction with the encoder and decoder, assigns a permutation matrix to each input graph. This matrix aligns the node order of the input graph with that of its reconstruction, ensuring consistent representation across different inputs.
Evolving instances with CMA-ES
We aim to search hard graph instances for a maximum cut solver against another solver. Hence, we propose using the following objective function:
(1) $\textit{p}_{\text{B}}(\text{I})/\textit{p}_{\text{A}}(\text{I}),$
where $A$ is the baseline/reference algorithm and $B$ is the one of interest. In this paper, we try to answer two questions: is there a cluster/family of graphs (1) hard to solve by RQAOA w.r.t. the classical GW method; (2) vice versa? For those two questions, we maximize the objective functions, $\textit{p}_{\text{GW}}(\text{I})/\textit{p}_{\text{RQAOA}}$ and $\textit{p}_{\text{RQAOA}}(\text{I})/\textit{p}_{\text{GW}}$, respectively, in the latent space of the already-trained PIGVAE model. We utilize the widely applied CMA-ES for this maximization task. For each latent point sampled by the CMA-ES, we first decode the latent point to a graph instance $I$ and then execute RQAOA and GW algorithms on instance $I$ 100 times due to their stochasticity. We take the median value of each algorithm to compute the above ratio.
Importantly, a randomly sampled latent point may be decoded into a disconnected graph. We might initiate the CMA-ES with connected graphs. Still, as the search progresses, CMA-ES may sample disconnected graphs again. To resolve this issue, we propose the following strategy: for any latent point that corresponds to a disconnected graph, we solve the maximum cut problems supported on each connected component and then take the sum of the number of cuts from all components. For RQAOA, if the number of nodes of a component is smaller than a preset value $\textit{n}_{c}$, we solve the maximum cut problem on this component by a brute force method; Otherwise, we apply RQAOA on it. In the recursion step of RQAOA, disconnectedness can be created even if we start with a connected graph. In this case, we apply the above strategy to the connected components. In the experiments, we test our approach with two sizes: 20 nodes and 100 nodes, and we set $\textit{n}_{c}$ to 10 and 20, respectively.
5. Experiments
5.1. Experimental setup
Figure 2. Second-largest eigenvalues of the normalized graph Laplacian for graph embeddings of 20-node (top) and 100-node (bottom) graphs. Nodes represent DBSCAN-identified clusters (Ester et al., 1996; Schubert et al., 2017), with edges indicating feature overlap or transitions between clusters. Node colors reflect the second-largest eigenvalue, offering insights into graph connectivity and structural complexity.
[…]
Second-largest eigenvalues of the normalized graph Laplacian for graph embeddings of 20-node (top) and 100-node (bottom) graphs.
GAE training
We sampled the training graph instance u.a.r. from six random graph models: Erdős–Rényi (Erdos and Rényi, 1960), Watts-Strogatz small-world (Watts and Strogatz, 1998), Newman-Watts-Strogatz small-world (Newman and Watts, 1999), Random Regular (Steger and Wormald, 1999), Barabási–Albert (Barabási and Albert, 1999), and Dual-Barabási–Albert graphs (Moshiri, 2018). We also took the parameter setting of those graph models and hyperparameters of the PIGVAE model from (Winter et al., 2021b). In each training epoch, we generate 6 000 graph instances from the above random models. We train the PIGVAE model with maximally 3 000 epochs and a batch size of 128. An early stopping mechanism is implemented to terminate training if the validation loss shows no improvement over the past 60 epochs. To assess our approach on different graph sizes, we train two separate models: one for 20-node instances and the other for 100-node instances.
CMA-ES setup
The CMA-ES algorithm is configured with a population size of 64, leveraging parallel evaluations to enhance computational efficiency. To further optimize performance, a restart mechanism is implemented, allowing up to 10 restarts, with each limited to a maximum of 2,000 function evaluations. The algorithm is based on the standard implementation provided by the pycma library (Hansen et al., 2019).
Reproducibility
The full implementation is available in our Zenodo repository (Authors, s), along with supplementary materials for reproducibility and reference.
5.2. Results
5.2.1. Evaluation of the GAE
During the training phase, we randomly save 60 graphs out of 6,000 in each epoch, accumulating a total of 6,000 graphs for in-depth evaluation. Our initial examination focuses on the 64-dimensional latent space structure, utilizing embeddings generated from these 6,000 graphs. Initially, we compute several key features, as detailed in Section 7.1 (see the supplementary material (Authors, s)), prioritizing those that significantly influence the complexity and difficulty of the maximum cut problem. Research indicates that certain graph characteristics, especially eigenvalues, provide substantial insight into a graph’s connectivity and complexity concerning the maximum cut problem (Dunning et al., 2018; Moussa et al., 2020). Figure 2 presents the second-largest eigenvalues of the normalized graph Laplacian for 20-node and 100-node graphs. Nodes represent graph embedding clusters identified via DBSCAN clustering (Ester et al., 1996; Schubert et al., 2017), with edges indicating feature overlap or transitions between clusters. Node color reflects the second-largest eigenvalue, revealing graph connectivity and structural complexity. The visualization demonstrates the smooth consistency of embeddings, validating the well-trained GAE’s suitability for generating new graphs for further analysis.
We continue to evaluate our trained GAE through a link prediction task, leveraging learned embeddings to capture graph topology. These embeddings facilitate adjacency matrix reconstruction, enabling predictions of potential node connections. Performance is measured using the macro area under the receiver operating characteristics curve (ROC-AUC) (0 to 1 scale, where 0.5 indicates random guessing and 1 denotes perfect prediction). Our initial assessment focuses on a GAE trained on 10-node Erdős–Rényi graphs (50% probability) using a single A100 GPU with a batch size of 32. This setup yields a macro ROC-AUC of $0.933\pm 0.004$ across 6,000 graphs training graphs. Building on this promising result, we extend our analysis to larger graphs containing all graph models, obtaining macro ROC-AUC scores of $0.709\pm 0.001$ for 20-node graphs and $0.611\pm 0.001$ for 100-node graphs, using the same number of training samples.
5.2.2. Feature analysis through parallel coordinate plot
(a) 20-node hard maximum cut instances.
(b) 100-node hard maximum cut instances.
Figure 3. Computed features of generated 20-node (top) and 100-node (bottom) hard maximum cut instances. Each pair of charts highlights cases where either the RQAOA or GW algorithm outperforms the other, with the color bar representing ratio values. Lower values indicate instances where one algorithm significantly outperforms the other.
[…]
Computed features of generated 20-node (top) and 100-node (bottom) hard maximum cut instances.
To ensure diversity, each CMA-ES run begins from different latent points within the latent space. For 20-node maximum cut instances, we generate 1,519 instances where RQAOA outperformed GW and 1,231 instances where GW exceeded the performance of RQAOA. Similarly, for 100-node instances, both algorithms outperform each other in 850 cases. To further analyze these instances, we examine various features (see section 7.1 in the supplementary material (Authors, s)). Comparative results are visualized in Figure 3(a) for 20-node graphs and Figure 3(b) for 100-node graphs.
Figure 3(a) presents feature comparisons for maximum cut instances where RQAOA and GW exhibit differing performance. The top subplot highlights cases where RQAOA outperforms GW, with darker shades indicating greater superiority. The bottom subplot shows instances favoring GW, revealing that lower performance ratios (indicating stronger performance) appear more frequently when GW surpasses RQAOA. RQAOA-dominant instances exhibit a narrow performance ratio range (0.952–0.979), whereas GW-favored cases span a wider range (0.720–0.990). A key distinction emerges in LOG_SMALLESTEIGVAL, where RQAOA-preferred graphs fall below -20, suggesting structural isolation. A representative case, featuring a mostly connected graph with one isolated node, illustrates this trend. Running 100 trials per algorithm on this instance, RQAOA outperforms GW in only 4 trials, with identical results in the rest, implying that GW’s inherent randomness may influence its comparative performance. Figure 3(b) reveals a similar trend for 100-node graphs, indicating no significant differences in computed features compared to the 20-node case.
5.2.3. Feature analysis with machine learning pipeline
Important score Feature
$0.1471\pm 0.0059$ EXPECTED_COSTGW_OVER_SDP_COST
$0.1370\pm 0.0084$ SPECTRAL_GAP
$0.0769\pm 0.0057$ PERCENT_CLOSE1_LOWER_TRIANGULAR
$0.0684\pm 0.0031$ PERCENT_POSITIVE_LOWER_TRIANGULAR
$0.0665\pm 0.0073$ LOG_SECONDLARGESTEIGVAL
$0.0528\pm 0.0042$ LOG_LARGESTEIGVAL
$0.0527\pm 0.0042$ DENSITY
$0.0521\pm 0.0056$ LOG_NUM_EDGES
$0.0508\pm 0.0051$ PERCENT_CLOSE3_LOWER_TRIANGULAR
$0.0434\pm 0.0068$ LOGRATIO_EDGETONODES
$0.0431\pm 0.0040$ NORM_MIS
$0.0413\pm 0.0036$ GRAPH_ASSORTATIVITY
$0.0408\pm 0.0042$ TRANSITIVITY
$0.0378\pm 0.0027$ STD_COSTGW_OVER_SDP_COST
$0.0369\pm 0.0049$ PERCENT_CUT
$0.0310\pm 0.0052$ LOG_SMALLESTEIGVAL
$0.0215\pm 0.0041$ LOG_CHROMATIC_NUM
$00.0000\pm 0.0$ 0.0000 GIRTH
Table 1. Important features with the classifier built on 20-node maximum cut instances.
Important score Feature
$0.3749\pm 0.0104$ EXPECTED_COSTGW_OVER_SDP_COST
$0.1588\pm 0.0196$ LOG_SMALLESTEIGVAL
$0.1281\pm 0.0052$ PERCENT_CLOSE1_LOWER_TRIANGULAR
$0.0648\pm 0.0081$ PERCENT_CUT
$0.0618\pm 0.0049$ PERCENT_CLOSE3_LOWER_TRIANGULAR
$0.0454\pm 0.0071$ DENSITY
$0.0322\pm 0.0054$ GRAPH_ASSORTATIVITY
$0.0239\pm 0.0049$ LOG_LARGESTEIGVAL
$0.0199\pm 0.0085$ LOG_NUM_EDGES
$0.0165\pm 0.0037$ NORM_MIS
$0.0153\pm 0.0036$ TRANSITIVITY
$0.0143\pm 0.0026$ LOG_SECONDLARGESTEIGVAL
$0.0126\pm 0.0061$ LOGRATIO_EDGETONODES
$0.0116\pm 0.0026$ LOG_CHROMATIC_NUM
$0.0096\pm 0.0030$ SPECTRAL_GAP
$0.0061\pm 0.0032$ PERCENT_POSITIVE_LOWER_TRIANGULAR
$0.0043\pm 0.0025$ STD_COSTGW_OVER_SDP_COST
$00.0000\pm 0.0$ 0.0000 GIRTH
Table 2. Important features with the classifier built on 100-node maximum cut instances.
To analyze the GW algorithm’s role in generating hard maximum cut instances, we expand our feature set based on prior research (Moussa et al., 2020). Incorporating six additional GW-related features (see Section 7.2 in the supplementary material (Authors, s)), we increase our total to 18 features within our machine learning pipeline. NP-hard features like the domination number are considered but excluded due to their computational complexity and minimal impact on model accuracy. For classification, we use the TPOP library (Olson et al., 2016), an AutoML tool for optimizing machine learning pipelines. Instances, where RQAOA outperforms GW, are labeled 0, while GW-dominant cases are labeled 1. We split the dataset 70:30 for training and testing. Due to class imbalance, particularly in 20-node experiments, we optimize for balanced accuracy (Brodersen et al., 2010), while standard accuracy (Pedregosa et al., 2011) is used for the balanced 100-node dataset. Using TPOP, we conduct 10-fold cross-validation over 100 generations, evaluating 60 candidate solutions per generation.
In our 20-node experiments, we achieve an outstanding average balanced accuracy of $0.9988\pm 0.0012$ across ten runs, each with a unique random data split. This result is obtained using an ExtraTreesClassifier (Geurts et al., 2006) (parameters detailed in section 7.3). Key predictive features include EXPECTED_COSTGW_OVER_SDP_COST and SPECTRAL_GAP, where the former evaluates GW’s effectiveness, and the latter reflects graph spectral properties as shown in Table 1. The relevance of the GW algorithm’s output in our analysis aligns with our focus on instances that are particularly challenging for maximum cut problems, influenced by the GW performance. The importance of eigenvalues also plays a crucial role, as seen in LOG_SECONDLARGESTEIGVAL and LOG_LARGESTEIGVAL, while GIRTH (shortest cycle length) contributes little to classifier performance. Further, the partial dependence plot visualizes the impact of two top features as shown in Figure 4(a). This plot demonstrates how these features influence the probability of predicting class 1, which in this study corresponds to instances where GW outperforms RQAOA. The plot reveals a clear trend: as the feature value changes, the probability of an instance being classified as class 1 increases, indicating a positive correlation between these features and the probability of challenging GW.
(a) Partial dependence plot showing the impact of the top two features on 20-node maximum cut instances.
(b) Partial dependence plot showing the impact of the top two features on 100-node maximum cut instances.
Figure 4. Partial dependence plots showing the classifier’s response to the top two features of 20-node (top) and 100-node (bottom) maximum cut instances. The contour lines represent different probability levels, providing insight into feature influence.
[…]
Partial dependence plots illustrating the classifier’s behavior based on the top two features of 20-node (top) and 100-node (bottom) maximum cut instances.
In the analysis of 100-node instances, we achieve perfect accuracy, recorded as $1.0000\pm 0.0000$, using an ExtraTreesClassifier. The parameters for this classifier are outlined in section 7.3. As indicated in Table 2, the feature EXPECTED_COSTGW_OVER_SDP_COST stands out as the most significant, showing a much higher influence compared to other features. This finding is in line with our earlier results from the 20-node experiments, where four out of the five most influential features are similarly linked to the performance metrics of the GW algorithm. Consistent with previous observations, the feature GIRTH, which do not impact the classifier’s decisions. The influence of the primary features is visualized in Figure 4(b) through a partial dependence plot. It illustrates that as the values of EXPECTED_COSTGW_OVER_SDP_COST and LOG_SMALLESTEIGVAL increase, there is a corresponding rise in the probability that an instance will perform better under GW compared to RQAOA.
To assess the significance of features, we conduct a permutation important analysis. This involved randomly shuffling the values within columns of our dataset, focusing particularly on the top two features identified in both the 20-node and 100-node experiments. For each feature, we generate 10 different permutations of our original dataset. We then evaluate the impact of these permutations on the model’s performance, which is initially trained on the unaltered dataset, using the same 10 seeds as the original setup. For the 20-node experiments, the permutation of the feature EXPECTED_COSTGW_OVER_SDP_COST results in a slight decrease in performance, with metrics changing by a $-0.0006\pm 0.0018$. Interstingly, shuffling the SPECTRAL_GAP feature leads to a minor performance increase, indicated by $0.0001\pm 0.0017$. When both features are permuted simultaneously, the combined effect causes a more noticeable performance decrease of $0.0009\pm 0.00177$. In the 100-node experiments, permutating the EXPECTED_COSTGW_OVER_SDP_COST feature lead to a more pronounced decline in performance, measured at $-0.0067\pm 0.0048$. Permutation of the feature LOG_SMALLEST
EIGVAL results in a smaller drop in performance, recorded as $-0.0004$
±$\pm 0.0008$ on performance due to the. Permuting both features together causes a decrease in performance of $-0.0012\pm 0.0013$. Overall, we can conclude that the performance shifts upon permuting a single feature or two top features slightly affect the model’s performance.
To determine whether classifier performance could be sustained without the top two most informative features identified in our 20-node and 100-node experiments, we re-initiate the entire automated machine learning process excluding these features. In the 20-node experiments, this exclusion results in a minor decrease in performance, showing a decline of $0.0015\pm 0.0022$ from the original baseline. Interestingly, in the 100-node experiments, performance remained unchanged, suggesting that the other features in the dataset still carry significant predictive power and are capable of compensating for the absence of the top two features. As indicated in the study (Moussa et al., 2020), two features associated with GW performance are identified as the most significant factors for building the classifiers. Motivated to further explore this, we test whether classifiers built solely on GW-related features could maintain similar levels of performance. The results show a substantial drop in the 20-node experiment with accuracy falling to $0.9250\pm 0.1541$, indicating that GW-related features alone are not sufficient at this scale. However, the 100-node experiment continues to achieve perfect accuracy of $1.0000\pm 0.0000$. These findings suggest that while the GW-related features are insufficient on their own for the 20-node experiments, they provide substantial support in building the classifier with our 100-node instances.
5.2.4. RQAOA as a high-performing heuristic with machine leaning pipeline
(a) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 20-node maximum cut instances, based on permutation importance.
(b) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 100-node maximum cut instances, based on permutation importance.
Figure 5. Partial dependence plots illustrating the classifier’s behavior in identifying significant performance gains with RQAOA for 20-node (top) and 100-node (bottom) maximum cut instances. The plots highlight the influence of the top two features in distinguishing performance improvements, as determined by permutation importance.
[…]
Partial dependence plots illustrating the classifier’s behavior in identifying significant performance gains with RQAOA for 20-node (top) and 100-node (bottom) maximum cut instances. The plots highlight the influence of the top two features in distinguishing performance improvements, as determined by permutation importance.
Important score Feature
$0.0665\pm 0.0641$ STD_COSTGW_OVER_SDP_COST
$0.0361\pm 0.0382$ DENSITY
$0.0355\pm 0.0389$ NORM_MIS
$0.0352\pm 0.0383$ GRAPH_ASSORTATIVITY
$0.0347\pm 0.0372$ LOG_LARGESTEIGVAL
$0.0345\pm 0.0376$ LOG_CHROMATIC_NUM
$0.0331\pm 0.0344$ TRANSITIVITY
$0.0319\pm 0.0201$ SPECTRAL_GAP
$0.0180\pm 0.0191$ LOG_SECONDLARGESTEIGVAL
$0.0082\pm 0.0051$ LOG_SMALLESTEIGVAL
$0.0031\pm 0.0034$ PERCENT_POSITIVE_LOWER_TRIANGULAR
$0.0031\pm 0.0039$ LOG_NUM_EDGES
$0.0028\pm 0.0049$ PERCENT_CUT
$0.0012\pm 0.0025$ LOGRATIO_EDGETONODES
$0.0010\pm 0.0048$ PERCENT_CLOSE1_LOWER_TRIANGULAR
$0.0002\pm 0.0051$ PERCENT_CLOSE3_LOWER_TRIANGULAR
$00.0000\pm 0.0$ 0.0000 GIRTH
$-0.0002\pm 0.0029$ EXPECTED_COSTGW_OVER_SDP_COST
Table 3. Important features with the classifier built on 20-node maximum cut instances demonstrate RQAOA’s high performance, underscored by its significant permutation importance.
Important score Feature
$0.0243\pm 0.0379$ EXPECTED_COSTGW_OVER_SDP_COST
$0.0166\pm 0.0339$ DENSITY
$0.0114\pm 0.0057$ LOGRATIO_EDGETONODES
$0.0066\pm 0.0110$ NORM_MIS
$0.0036\pm 0.0106$ LOG_SECONDLARGESTEIGVAL
$0.0028\pm 0.0047$ LOG_NUM_EDGES
$0.0023\pm 0.0048$ LOG_LARGESTEIGVAL
$0.0011\pm 0.0020$ PERCENT_POSITIVE_LOWER_TRIANGULAR
$0.0010\pm 0.0052$ LOG_CHROMATIC_NUM
$0.0003\pm 0.0027$ STD_COSTGW_OVER_SDP_COST
$0.0003\pm 0.0004$ SPECTRAL_GAP
$0.0001\pm 0.0002$ TRANSITIVITY
$00.0000\pm 0.0$ 0.0000 GIRTH
$-0.0001\pm 0.0004$ PERCENT_CLOSE1_LOWER_TRIANGULAR
$-0.0001\pm 0.0003$ PERCENT_CLOSE3_LOWER_TRIANGULAR
$-0.0002\pm 0.0004$ PERCENT_CUT
$-0.0002\pm 0.0008$ GRAPH_ASSORTATIVITY
$-0.0153\pm 0.0118$ LOG_SMALLESTEIGVAL
Table 4. Important features with the classifier built on 100-node maximum cut instances demonstrate RQAOA’s high performance, underscored by its significant permutation importance.
After successfully applying our evolutionary approach to evolve hard maximum cut instances, a new question emerges: whether RQAOA consistently performs well? For our purposes, we seek to establish if RQAOA can achieve a performance level above 0.96, with a minimum advantage of 0.04 over the GW algorithm. In this study, we reassess the same dataset with a modified labelling strategy. Instances are labelled as 1 when the ratio of GW to RQAOA is 0.96 or less, indicating at least a 4% improvement over the GW algorithm. All other instances are labelled as 0. We develop classifiers on both 20-node and 100-node instances that incorporate all the features previously discussed with achieved a balanced accuracy of $0.9826\pm 0.0042$ on 20-node experiments and a $0.8343\pm 0.0303$ in the 100-node experiment, maintaining the same experimental and analysis procedures as before. The models, generated using the TPOP library are detailed further in section 7.3. Figure 5 shows the partial dependent plot for the top two features in both 20-node and 100-node cases, respectively. The calculation of feature importance scores is based on permutation importance (Pedregosa et al., 2011). The analysis of these plots reveals a clear pattern: low values in GW-related features are strongly associated with increased predicted probabilities of RQAOA achieving outstanding performance. Moreover, the significance of the DENSITY feature in influencing RQAOA’s performance is evident. This analysis not only shows the impact of GW-related cost features but also emphasizes the critical role of DENSITY in determining the effectiveness of the RQAOA algorithm.
Table 3 and Table 4 showcase the key features identified by the classifier for 20-node and 100-node maximum cut instances, respectively. Highlighted in these tables, the features emphasize the robust performance of RQAOA, validating its effectiveness as a high-performing heuristic through significant permutation importance.
We evaluate the performance of RQAOA against different thresholds to better understand our generated instances. Initially, thresholds of 0.95 and 0.98 are also considered; however, they are discarded because there are different situations related to 20-node or 100-node instances that no instances meet the criteria to be labelled as 1. At a threshold of 0.97, the average balanced accuracy achieved is $0.9795\pm 0.0045$ for 20-node instances, which is comparable to the performance at a threshold of 0.96. LOG_CHROMATIC_NUM and GRAPH_ASSORTATIVITY are identified as the most significant features according to permutation importance calculations. For 100-node instances, the performance improved notably at the 0.97 threshold, achieving an average balanced accuracy of $0.9214\pm 0.0049$. This represents a substantial enhancement over the results obtained with the 0.96 threshold. The most influential features for this larger node size are identified as PERCENT_CLOSE3_LOWER_TRIANGULAR and GRAPH_ASSORTATIVITY. These results demonstrate that the model’s effectiveness varies significantly with different threshold settings, with each threshold providing unique insights, especially in scenarios involving larger graph instances. This responsiveness underscores the need to thoughtfully choose performance benchmarks that align with the unique attributes and requirements of the dataset.
6. Conclusion
Given the current literature does not provide detailed guidelines for constructing hard maximum cut instances, our research aims to address this gap. We focus on developing instances that highlight the distinctions between the quantum method RQAOA and the classical GW algorithm by effectively identifying maximum cut instances that pose significant challenges or are notably manageable for the RQAOA compared to the performance of the GW algorithm. Specifically, our work extends beyond situations where RQAOA merely surpasses GW, exploring instances where RQAOA achieves exceptionally high-quality outcomes and significantly outperforms GW. We employ an evolutionary approach using the classical CMA-ES algorithm, enhanced with a fitness function specifically tailored within the latent space of a well-trained Graph Autoencoder.
Drawing on precedents that applied machine learning to analyze the performance of quantum methods (Melnikov et al., 2019; Moussa et al., 2020), we adapt a comprehensive machine-learning pipeline. This approach examines a wide array of features, from those common to regular graphs and the spectrum of the Laplacian matrix to those associated with the relaxed solution of the semi-definite program in GW. Our feature analysis of the generated instances has further substantiated that RQAOA can offer substantial benefits over the GW algorithm.
In predicting these advantages, the machine learning model we developed strong results across both 20-node and 100-node instances. It confirms the necessity of both categories of features in analyzing instances where RQAOA significantly outperforms GW, reinforcing the comprehensive nature of the analytical framework. Our findings highlight EXPECTED_COSTGW_OVER_SDP_COST as the most influential feature when RQAOA is not considered a high-performing heuristic. This trend remains consistent in 100-node maximum cut instances when we consider RQAOA as a high-performing heuristic but does not hold for 20-node instances. Conversely, DENSITY plays a key role when RQAOA functions as a high-performing heuristic. Beyond these primary features, various other features also contribute to the machine learning classifiers, but GIRTH is found to have no impact in either experimental setting.
Our study offers a fresh perspective on graph generation within network science, encouraging further exploration and advancements in this area. Future work can be explored from various perspectives. Our current experiments focus on graphs with 20 and 100 nodes, but we are actively working to scale up to larger instances, such as 400-node graphs. A major challenge in this expansion is the scalability of the PIGVAE model, which was originally developed for smaller graphs and faces significant memory constraints when applied to larger ones. Running the model on 400-node graphs leads to severe computational bottlenecks. To address this, we have reduced the model’s size, but this comes at the potential cost of training performance. Moving forward, we aim to explore alternative strategies to improve scalability, enabling the model to efficiently process graphs with thousands of nodes while maintaining its effectiveness. Additionally, as part of our future work, we aim to integrate additional heuristics beyond GW to enhance the depth of our comparative analysis.